OpenStudy (anonymous):

Lab is to find the radius and height of a right circular cylinder where the surface area and volume are the same number. Is there a solution? Is there more than one?

OpenStudy (campbell_st):

well the surface area is \[SA = 2\pi r^2 + 2\pi r h\] and the volume is \[V = \pi r^2 h\] so if they are equal to the same number \[\pi r^2 h = 2\pi r^2 + \pi r h\] and then work from there..

OpenStudy (anonymous):

what should I choose for the number? any?

OpenStudy (campbell_st):

so if you have \[\pi r^2 h = 2 \pi r^2 + 2 \pi r h\] divide every term by \[\pi r\] then \[rh = 2r + 2h\] make h the subject \[h( r -2) = 2r\] then \[h = \frac{2r}{r + 2}\]

OpenStudy (campbell_st):

that's my best guess at this question... so r > 0

OpenStudy (campbell_st):

so as r gets large it looks like h gets closer to 2...

OpenStudy (anonymous):

ok thanks